# Approaching the thermodynamic limit of a bounded one-component plasma

### v1

@@ III.2 Simulation results
Figure 1 shows typical ion density distributions and the radial distribution functions obtained in our simulations. Largely, in the BOCP core region, the radial distribution functions are almost identical to those obtained for OCP in various MC simulations starting from [40] (Fig. 1b). Namely, a dip in $f_{p}(r)$ at small distances for $\Gamma=0.1$ is indicative of repulsion, the radial distribution function exhibits two maxima indicating a smooth transfer from the gaslike to fluid state at $\Gamma=30$, and several maxima emerge at $\Gamma=200$, which is somewhat below the melting point for the size $N=5000$. The maxima at larger $r$ are due to the shells formed closer to the BOCP boundary (two shells can be seen already at $\Gamma=30$). Note that melting of shells is observed at $\Gamma=60\mbox{--}90$ well below the melting point of a central part of the Coulomb clusters [64]. In Fig. 1a, the ion density distribution $f_{c}(r)$ demonstrates qualitatively the same regularities: in most of the BOCP volume, the ions are homogeneously distributed at $\Gamma=0.1$ and 30, although two shells are visible for $\Gamma=30$. However, the curve for $\Gamma=200$ is more indicative of the shell structure than in Fig. 1b because in the latter, departure of the ion, for which $f_{p}(r)$ is determined, from the BOCP center smooths the peaks of shells considerably. Instead, as follows from Fig. 1a, shells fill a significant part of the BOCP volume at high $\Gamma$. As is seen in Fig. 1c, both $f_{p}(r)$ and $f_{c}(r)$ show behavior with irregular tooth-shaped peaks typical for a solid.
Figure 1: Ion density distribution $f_{c}(r)$ (8) and the radial distribution function $f_{p}(r)$ (7) of BOCP at different $\Gamma$: (a) and (b) $\Gamma=0.1$, 30, and 200 (red, green, and blue lines, respectively), $N=5000$; (c) $f_{c}(r)$ (red line) and $f_{p}(r)$ (blue line) at $\Gamma=1000$ and $N=14000$.
@@
The size dependence of $u(\Gamma,N)$ obtained from MD simulation is illustrated by Fig. 2a, which testifies the accuracy of its approximation by the function (19). Indeed, the actual deviation of this approximation from the obtained MD results for all BOCP sizes, e.g., for $\Gamma=30$, varies from $4\times 10^{-8}$ to $8\times 10^{-6}$. The accuracy of MD results determined by a standard statistical method depending on the sample size is of the same order of magnitude and principally it can be further decreased by the increase in $N_{\tau}$. However, the difference in $u(\Gamma,N)$ obtained for two similar runs initialized with different random seeds is orders of magnitude greater, which means that a real simulation error is of physical nature. Thus, this difference may be due to the low-frequency non-vanishing wing of the phonon spectrum when the inverse phonon frequency is of the same order as $\tau N_{\tau}$, or it may arise from different structures of the polycrystal formed at $\Gamma\gg 1$. Thus, for each $\Gamma$ and $N$, the simulation was repeated twice; $u(\Gamma,N)$ was then estimated as the average of these two results while the accuracy was set to the modulus of their difference. For, $\Gamma=30$, such accuracy does not exceed $4\times 10^{-5}$, which is noticeably greater than the accuracy of data approximation. The same procedure was applied for all determined size-dependent quantities. It is this accuracy that is shown by error bars in Figs. 2, 4, and 68. Note that in most figures, the error bar width is less than the dot marker size.
@@
Albeit the function (19) approximates $u(\Gamma,N)$ with a high accuracy, the size dependence is very weak at great $N$. This makes problematic extrapolation of $u(\Gamma,N)$ to the thermodynamic limit. In other terms, it is necessary to estimate the accuracy of $u_{\infty}(\Gamma)$ obtained from such extrapolation. To this end, from the set of data for the same $\Gamma$, we selected at random two triplets of unmatched BOCP sizes $N$ and fitted (19) to these triplets. Modulus of the half-difference between two values of $u_{\infty}(\Gamma)$ obtained in such a way is associated with the accuracy of its determination in Figs. 3 and 5. As is seen in Table 2, the overall relative accuracy of extrapolated quantities is of the order of 0.1%, which is greater than the errors for individual values but still reasonably good. Figure 2b illustrates an extremely slow convergence of $u(\Gamma,N)$ to $u_{\infty}(\Gamma)$. As is seen, the accuracy of 1% is reached already for $N=45000$, which is close to the size used in this work, while the overall approximation accuracy of 0.1% is attained only for $N=10^{34}$, i.e., for a large body.
@@
Table 2: BOCP parameters in the thermodynamic limit. Listed are the total potential energy $u_{\infty}$ and its fit (38), (39); the excess energy of interionic interaction $u_{p\mathrm{ex}}$, the excess energy of interaction between the ions and background $u_{b\mathrm{ex}}$, and the shifted ion compressibility factor $Z_{i}-1$ calculated from the excess energies (16) and virial (24) (for $u_{p\mathrm{ex}}$ and $u_{b\mathrm{ex}}$, the results for $N=21200$ are listed).
$\Gamma$} $u_{\infty}$ $u_{\mathrm{fit}}$ $u_{p\mathrm{ex}}$ $u_{b\mathrm{ex}}$ $Z_{i}-1$ (energy) $Z_{i}-1$ (virial)
@@
Figure 3: Thermodynamic limit of the BOCP electrostatic energy compared with the results of MC simulations for the OCP with PBC. MD results of this work (circles) and their approximation by the functions (38) and (39) (solid line), MC results by Brush et al. [40] (squares), Hansen [41] (diamonds), Caillol and Gilles [46] (stars), Demyanov and Levashov [47] (triangles). Approximation of the results available in studies by Slattery et al. [43] and Caillol (MC within hyperspherical boundary conditions) [45] are shown by dotted and dashed-dotted lines, respectively. Dashed-dotted-dotted line indicates the Debye–Hückel approximation and dashed line, the Madelung constant for bcc lattice. Inset shows a magnified fragment of this figure.
The total electrostatic energy $u_{\infty}(\Gamma)$ determined from MD simulation is listed in Table 2. These data were fitted by the function borrowed from [45]:
$\Gamma{u_{\mathrm{fit}}}=A+B\Gamma+C{\Gamma^{s}}+D{\Gamma^{-s}},\quad 0.1\leq\Gamma\leq 174,$ (38)
with A=0.8304212370A=-0.8304212370, B=0.8974963531B=-0.8974963531, $C=0.9500252063$, $D=0.19951975244$, and $s=0.2396352836$, and by the theoretical dependence (cf. (25) and (26))
@@
with $\beta_{0}=-0.8963837125$, $\beta_{1}=1.5311038492$, and $\beta_{2}=35.092640432$. Note that at $\Gamma\leq 176$, $\left|{{u_{\infty}}-{u_{\mathrm{fit}}}}\right|\sim{10^{-4}}$, which is much less than the error in $u_{\infty}$ definition.
@@
The parameters in Eq. (39) obtained from the fitting procedure prove to be close to those calculated in the harmonic lattice approximation: $\beta_{0}\approx u_{0}$ with the same accuracy of $\sim 10^{-3}$ as in Table 2 and $\beta_{1}\approx 3/2$. The first anharmonicity correction coefficient turns out to be of the same order of magnitude as calculated in study by Chugunov and Baiko [26] for the classical OCP, $\beta_{2}\approx 10.58$. Apparently, the accuracy of $\beta_{2}$ determination could be enhanced if more points were calculated in the interval $174<\Gamma\leq 1000$ (at present, only three were fitted), and it might bring our result closer to [26].
@@
Determined energies are compared with the results of different studies in Fig. 3. As is seen, they are in a good agreement with early data by Brush et al. [40] and Hansen [41], however, dispersion of these data are much greater than the accuracy attained in this work. Note that here, our error bars are smaller than the dot marker sizes. The results obtained later by Slattery et al. [43] are in perfect agreement with modern data from the works by Demyanov and Levashov [47], and by Caillol [45] and Caillol and Gilles [46] for $\Gamma\leq 1$, while at $\Gamma>1$, the data from [46] are situated higher than any other data. At $\Gamma\leq 0.1$, our results scarcely deviate from the Debye–Hückel approximation, while the data from [46] are noticeably higher. At $\Gamma>174$, our data reach asymptotically the bcc energy $u_{0}$, and in this region, they are in a very good agreement with [43] (Fig. 5). At the same time, for $0.1<\Gamma<174$, our energies are appreciably lower than other data, which can be clearly seen in the inset in Fig. 3. Here, a typical difference is about 0.5%, which is much higher than the declared accuracy of all results. Apparently, this difference may arise from the limitation on system microscopic states imposed by the PBC, which can be qualitatively treated as strong correlations between the particles in a supercell and its images. Since the characteristic decay length of the Ewald potential is of the order of the simulation cell length, the resulting long-range interaction may contribute noticeably to the total system energy. In contrast, BOCP does not have such limitation, and this leads the system to a more favorable state. However, at $\Gamma\to\infty$, the system occupies a single state, which results in convergence of [43] and our data.
@@
Based on Eqs. (38) and 39), we can approximate the internal energy in a wide range of $\Gamma$ in the form
@@
and the Helmholtz free energy, as
${f_{i\mathrm{ex}}}(\Gamma)$ $=\int{\displaystyle{{{\varepsilon_{i\mathrm{ex}}}(\Gamma)}\over\Gamma}}{\mkern 1.0mu}{\kern 1.0pt}d\Gamma$ (41)
@@
${\omega_{2}}(\Gamma)=-{\beta_{1}}/\Gamma-{\beta_{2}}/2{\Gamma^{2}}$, $\Delta{\omega_{1}}={\omega_{2}}({\Gamma_{\mathrm{m}}})-{\omega_{1}}({\Gamma_{\mathrm{m}}})$, and we take into account that $f_{i\mathrm{e}x}(\Gamma)$ is a continuous function at $\Gamma=\Gamma_{\mathrm{m}}$.
@@
A good parameter for investigation of melting is the thermal fraction of the Coulomb energy $\Gamma[u(\Gamma,N)-u_{0}]$ [43]. It is well known that, in contrast to the bulk system, phase transitions in finite-size systems are very diffuse, both in temperature and phase composition. As can be seen in Fig. 4, BOCP is not exclusion. In this case, melting can be associated with the Coulomb coupling parameter $\Gamma_{\mathrm{m}}$ corresponding to the maximum of $\Gamma_{\mathrm{m}}[u(\Gamma_{\mathrm{m}},N)-u_{0}]$. This maximum is shifted toward lower values with the increase in $N$. Inset shows fitting $\Gamma_{\mathrm{m}}$ with the exponential decay function $\Gamma_{\mathrm{m}}(N)=174.0+647.995\exp(-N/1729.886)$, from which the estimate for the bulk melting parameter is $\Gamma_{\mathrm{m}}(\infty)=174\pm 2$. The accuracy of 2 corresponds to the error of maximum resolution for the size-dependent function $\Gamma_{\mathrm{m}}[u(\Gamma_{\mathrm{m}},N)-u_{0}]$ due to the error in $u(\Gamma,N)$. Obtained value correlates with that obtained in [43] ($\Gamma_{\mathrm{m}}=178$) and coincides with the results by Ichimaru [2], Dubin and O’Neil [14] and Khrapak and Khrapak [39]. The absence of appreciable metastable region in the neighborhood of $\Gamma_{\mathrm{m}}$ is noteworthy. It means that the fluid–solid transition is very fast, at least, its characteristic time is less than $N_{\tau}$. This can be connected with the solid-like shells near the BOCP boundary, which initiate crystallization propagating to the BOCP center.
@@
Figure 5: Thermal fraction of the electrostatic energy in the thermodynamic limit as a function of $\Gamma$. This work: MD results (dots) and their approximation by the function (38) (dashed line) and (39) (dashed-dotted line). Solid line indicates calculations based on the harmonic lattice approximation [26] and dotted line, MC results [43].
@@
Another way to estimate $\Gamma_{\mathrm{m}}$ is based on the parameter $\Gamma(u_{\infty}-u_{0})$ for the bulk fluid and solid (Fig. 5). Here, in contrast to Fig. 4, the dependence of this parameter on $\Gamma$ forms a cliff at $\Gamma_{\mathrm{m}}=174\pm 2$, where again, the estimation for accuracy follows from the error in $u_{\infty}(\Gamma,N)$. Note that at $190\leq\Gamma\leq 210$, the dots are situated well below the data for $\Gamma\geq 300$. This is a consequence of a bad applicability of the principal size dependence (19) for $N>10000$, so that the data in Table 2 were obtained only for the smaller sizes. Very large errors of $u_{\infty}$ definition in this range of $\Gamma$ are indicative of enormous fluctuations of the BOCP structure at $\Gamma\approx\Gamma_{\mathrm{m}}$. Here, $N_{\tau}$’s limited by the computational resources and listed in Table 1 seem to be insufficient to reach equilibration for such sizes. However, one can testify that at high $\Gamma$, our data agree qualitatively with the analytical results by Chugunov and Baiko [26]. Note that, as for a finite size BOCP, we have found no metastable region. In addition, it is worth mentioning that the excess free energy at the melting point (28) amounts to 0.0092002, which proves to be an order of magnitude less than what follows from the estimate of free energy in [43].
@@
Figure 6: Size dependences of the excess interatomic electrostatic energy $u_{p{\mathrm{e}x}}$ (circles) and the excess ion–background electrostatic energy $u_{b{\mathrm{e}x}}$ (squares) at $\Gamma=30$.
@@
Now turn to determination of the excess interatomic electrostatic energy and the excess ion–background electrostatic energy (Table 2 and Fig. 6). The size dependences of excess energies $u_{p\mathrm{e}x}(\Gamma,N)$ and $u_{b\mathrm{e}x}(\Gamma,N)$ (5) determined in our simulation reveal an overall trend to decrease with the increase in $N$. However, the decrease is so small that we failed to assign any functional dependence to it in the investigated range of system parameters. Rigorously speaking, we are not able to extrapolate the results to the thermodynamic limit due to an extremely narrow abscissa range in Fig. 6 and hence, confine ourselves to its estimate by taking these quantities for $N=21200$. Here, the corresponding errors were calculated in the same way as for $u(\Gamma,N)$. The resulting dependences $u_{p{\mathrm{e}x}}(\Gamma)$ and $u_{b{\mathrm{e}x}}(\Gamma)$ are shown in Fig. 7 and Table 2. Since at $182\leq\Gamma\leq 210$, BOCP of the sizes $N>10000$ were not simulated, the corresponding $u_{p{\mathrm{e}x}}$ and $u_{b{\mathrm{e}x}}$ are not listed in table. Behavior of these dependences is in agreement with the asymptotic analytical predictions (12) and (18), which is indicative of the fact that Fig. 7 represents a satisfactory estimate for the investigated quantities in the thermodynamic limit. One can see that as $\Gamma$ is decreased, $u_{b{\mathrm{e}x}}(\Gamma)$ vanishes more rapidly, which can be explained by the difference in $f_{p}(r)$ (7) and $f_{c}(r)$ (8). Indeed, at $\Gamma=0.1$, $f_{c}\approx 1$ (Fig. 1a), which vanishes the integral (10), whereas $f_{p}$ is noticeably different from zero at short distances (Fig. 1b) resulting in nonzero $u_{p{\mathrm{e}x}}(\Gamma)$ (9). Since (9) and (10) were written in the approximation of pair correlations for a homogeneous medium, it is clear that such approximation is ineffective for $u_{b{\mathrm{e}x}}(\Gamma)$, whose analytical calculation would require the inclusion of higher-order correlations. In this connection, a fast decrease of $u_{b{\mathrm{e}x}}(\Gamma)$ at $\Gamma>0.3$ is indicative of inadequacy of the pair correlations approximation, and therefore, $\Gamma=0.3$ can be regarded as an upper bound for the Debye–Hückel approximation or the weakly coupled regime. Notably, it follows from the recent analysis [55] of the reduced coefficients of self-diffusion, shear viscosity, and thermal conductivity of OCP based on the Debye–Hückel plus hole approach [51] that the binary collision approaches break down at $\Gamma>0.3$, when OCP is no more a weakly coupled system.
@@
At high $\Gamma$, $u_{p{\mathrm{e}x}}(\Gamma)$ reaches a solid state value already at $\Gamma=30$ while $u_{b{\mathrm{e}x}}(\Gamma)$, only at $\Gamma=650$, which is indicative of strong ionic correlations in this region. Interestingly, $u_{b{\mathrm{e}x}}(\Gamma)$ has a shallow minimum at $\Gamma\approx 10$, which coincides with a gas-to-liquid dynamical crossover in OCP discussed in works [21] and [54], so this minimum can serve as yet another definition of the Frenkel line in OCP.

### v2

@@ III.2 Simulation results
Figure 1 shows typical ion density distributions and the radial distribution functions obtained in our simulations. Largely, in the BOCP core region, the radial distribution functions are almost identical to those obtained for the OCP in various MC simulations starting from [51] (Fig. 1b). Namely, a dip in $f_{p}(r)$ at small distances for $\Gamma=0.1$ is indicative of repulsion; the radial distribution function exhibits two maxima indicating a smooth transfer from the gaslike to fluid state at $\Gamma=30$; and several maxima emerge at $\Gamma=200$, which is somewhat below the melting point for the size $N=5000$. The maxima at larger $r$ are due to the shells formed closer to the BOCP boundary (two shells can be seen already at $\Gamma=30$). Note that melting of shells is observed at $\Gamma=60\mbox{--}90$ well below the melting point of the central part of the Coulomb clusters [75]. In Fig. 1a, the ion density distribution $f_{c}(r)$ demonstrates qualitatively the same regularities: in most of the BOCP volume, the ions are homogeneously distributed at $\Gamma=0.1$ and 30, although two shells are visible for $\Gamma=30$. However, the curve for $\Gamma=200$ is more indicative of the shell structure than in Fig. 1b because in the latter, departure of the ion, for which $f_{p}(r)$ is determined, from the BOCP center smooths the peaks of shells considerably. Instead, as follows from Fig. 1a, shells fill a significant part of the BOCP volume at high $\Gamma$. As is seen in Fig. 1c, both $f_{p}(r)$ and $f_{c}(r)$ exhibit behavior with irregular tooth-shaped peaks typical for a solid. If we model the BOCP by a quasi-uniform core surrounded by shells then the radius of mesoscopic uniformity can be estimated as $N_{\mathrm{cr}}^{1/3}$, where $N_{\mathrm{cr}}$ is the number of ions forming a 3D crystallized core (not belonging to shells). As follows from the results of [73], for $N=5000$, $N_{\mathrm{cr}}^{1/3}\simeq 6.88$ at $\Gamma=210$ and $13.4$ at $500$, while the cluster radius is $R=17.1$ (see also Fig. 4 in [73]).
Figure 1: Ion distribution functions: (a) the ion density distribution $f_{c}(r)$ (8) and (b) the radial distribution function $f_{p}(r)$ (7) of BOCP at different $\Gamma$; (a) and (b) $\Gamma=0.1$, 30, and 200 (red, green, and blue lines, respectively), $N=5000$; (c) $f_{c}(r)$ (red line) and $f_{p}(r)$ (blue line) at $\Gamma=1000$ and $N=14000$. The BOCP radii are (a) 17.1, (b) 17.1, and (c) 24.1.
@@
The size dependence of $u(\Gamma,N)$ obtained from MD simulation is illustrated by Fig. 2a, which testifies the accuracy of its approximation by the function (19). Indeed, the actual deviation of this approximation from the obtained MD results for all BOCP sizes, e.g., for $\Gamma=30$, varies from $4\times 10^{-8}$ to $8\times 10^{-6}$. The accuracy of MD results determined by a standard statistical method depending on the sample size is of the same order of magnitude and in principle it can be further decreased by the increase in $N_{\tau}$. However, the difference in $u(\Gamma,N)$ obtained for two similar runs initialized with different random seeds is orders of magnitude greater, which means that a real simulation error is of physical nature. Thus, this difference may be due to the low-frequency non-vanishing wing of the phonon spectrum when the inverse phonon frequency is of the same order as $\tau N_{\tau}$, or it may arise from different structures of the polycrystal formed at $\Gamma\gg 1$. Thus, for each $\Gamma$ and NN, the simulation was repeated twice; $u(\Gamma,N)$ was then estimated as the average of these two results while the accuracy was set to the modulus of their difference. For $\Gamma=30$, such accuracy is not worse than $4\times 10^{-5}$, which is noticeably greater than the accuracy of data approximation. The same procedure was applied for all determined size-dependent quantities. It is this accuracy that is shown by error bars in Figs. 2, 4, and 68. Note that in most figures, the error bar width is less than the dot marker size.
@@
Albeit the function (19) approximates $u(\Gamma,N)$ with a high accuracy, the size dependence is very weak at great NN. This makes extrapolation of $u(\Gamma,N)$ to the thermodynamic limit problematic. In other terms, it is necessary to estimate the accuracy of $u_{\infty}(\Gamma)$ obtained from such extrapolation. To this end, from the set of data for the same $\Gamma$, we selected at random two triplets of unmatched BOCP sizes NN and fitted (19) to these triplets. Modulus of the half-difference between two values of $u_{\infty}(\Gamma)$ obtained in such a way is associated with the accuracy of its determination in Figs. 3 and 5. As is seen in Table 2, the overall relative accuracy of extrapolated quantities is of the order of 0.1%, which is greater than the errors for individual values but still reasonably good. Figure 2b illustrates an extremely slow convergence of $u(\Gamma,N)$ to $u_{\infty}(\Gamma)$. As is seen, the accuracy of 1% is reached already for $N=45000$, which is close to the size used in this work, while the overall approximation accuracy of 0.1% is attained only for $N=10^{34}$, i.e., for a large body.
@@
Table 2: BOCP parameters in the thermodynamic limit. Listed are the total potential energy $u_{\infty}$ and its fit $u_{\mathrm{fit}}$ (38), (39); the excess energy of interionic interaction $u_{p\mathrm{ex}}$, the excess energy of interaction between the ions and background $u_{b\mathrm{ex}}$ (listed are the results for $N=21200$), and the shifted ion compressibility factor $Z_{i}-1$ calculated from the excess energies (16) and virial (24).
$\Gamma$} $u_{\infty}$ $u_{\mathrm{fit}}$ $u_{p\mathrm{ex}}$ $2u_{b\mathrm{ex}}$ $Z_{i}-1$ (energy) $Z_{i}-1$ (virial)
@@
Figure 3: Thermodynamic limit of the BOCP electrostatic energy compared with the results of MC simulations for the OCP with PBC. Shown are the MD results of this work (circles) and their approximation by the functions (38) and (39) (solid line), MC results by Brush et al. [51] (squares), Hansen [52] (diamonds), Caillol and Gilles [57] (stars), Demyanov and Levashov [58] (triangles). Approximation of the results available in studies by Slattery et al. [54] and Caillol (MC within hyperspherical boundary conditions) [56] are shown by dotted and dashed-dotted lines, respectively. Dashed-dotted-dotted line indicates the Debye–Hückel approximation and dashed line, the Madelung constant for bcc lattice. Inset shows a magnified fragment of this figure.
The total electrostatic energy $u_{\infty}(\Gamma)$ determined from MD simulation is listed in Table 2. These data were fitted by the function taken from [56]
$\Gamma{u_{\mathrm{fit}}}=A+B\Gamma+C{\Gamma^{s}}+D{\Gamma^{-s}},\quad 0.1\leq\Gamma\leq 174$ (38)
with A=0.830421A=-0.830421, B=0.897496B=-0.897496, $C=0.950025$, $D=0.19952$, and $s=0.239635$, and by the theoretical dependence (cf. (25) and (26))
@@
with $\beta_{0}=-0.896383$, $\beta_{1}=1.5311$, and $\beta_{2}=35.0926$. Note that at $\Gamma\leq 176$, $\left|{{u_{\infty}}-{u_{\mathrm{fit}}}}\right|\sim{10^{-4}}$, which is much less than the error in $u_{\infty}$ determination.
@@
The parameters in Eq. (39) $\beta_{0}$ and $\beta_{1}$ obtained from the fitting procedure prove to be close to those calculated in the harmonic lattice approximation: $\beta_{0}\approx u_{0}$ with the same accuracy of $\sim 10^{-3}$ as in Table 2 and $\beta_{1}\approx 3/2$. The first anharmonicity correction coefficient turns out to be of the same order of magnitude as calculated in study by Chugunov and Baiko [34] for the classical OCP, where $\beta_{2}\approx 10.58$. Apparently, the accuracy of $\beta_{2}$ determination could have been improved if more points were calculated in the interval $174<\Gamma\leq 1000$ (at present, only three were fitted), and it might bring our result closer to [34]. Even better accuracy in $\beta_{0}$ is attained if we fit solely this parameter taking the theoretical values $\beta_{1}=1.5$ and $\beta_{2}=10.58$. Here, $\beta_{0}=-0.896204$, which is less than $u_{0}$ by only $2.8\times 10^{-4}$ . One can conclude that the energy of a cold system tends to $u_{0}$, which is in agreement with the result of work by Hasse [43].
@@
Determined energies are compared with the results of different studies in Fig. 3. As is seen, they are in a good agreement with early data by Brush et al. [51] and Hansen [52], however, dispersion of those data are much greater than the accuracy attained in this work. Note that here, our error bars are smaller than the dot marker sizes. The results obtained later by Slattery et al. [54] are in perfect agreement with modern data from the works by Demyanov and Levashov [58], and by Caillol [56] and Caillol and Gilles [57] for $\Gamma\geq 1$, while at $\Gamma<1$, the data from [57] are situated higher than any other data. At $\Gamma\leq 0.1$, our results scarcely deviate from the Debye–Hückel approximation, while the data from [57] are noticeably higher. At $\Gamma>174$, our data reach asymptotically the bcc energy $u_{0}$, and in this region, they are in a very good agreement with [54] (Fig. 5). At the same time, for $0.1<\Gamma<174$, our energies are appreciably lower than other data, which can be clearly seen in the inset in Fig. 3. Here, a typical difference is about 0.5%, which is much higher than the declared accuracy of all results. Apparently, this difference may arise from the limitation on system microscopic states imposed by the PBC, which can be qualitatively treated as strong correlations between the particles in a supercell and its images. Since the characteristic decay length of the Ewald potential is of the order of the simulation cell length, the resulting long-range interaction may contribute noticeably to the total system energy. In contrast, BOCP does not have such limitation, and this leads the system to a more favorable state. Nevertheless, at $\Gamma\to\infty$, the system occupies a single state, which results in closeness of our data and [54].
@@
Based on Eqs. (38) and (39), we can approximate the excess internal energy in a wide range of $\Gamma$ in the form
@@
and the excess Helmholtz free energy, as
${f_{\mathrm{ex}}}(\Gamma)$ $=\int{\displaystyle{{{\varepsilon_{\mathrm{ex}}}(\Gamma)}\over\Gamma}}{\mkern 1.0mu}{\kern 1.0pt}d\Gamma$ (41)
@@
${\omega_{2}}(\Gamma)=-{\beta_{1}}/\Gamma-{\beta_{2}}/2{\Gamma^{2}}$, $\Delta{\omega_{1}}={\omega_{2}}({\Gamma_{\mathrm{m}}})-{\omega_{1}}({\Gamma_{\mathrm{m}}})$, and we take into account that $f_{\mathrm{e}x}(\Gamma)$ is a continuous function at $\Gamma=\Gamma_{\mathrm{m}}$.
@@
A good parameter for investigation of melting is the thermal fraction of the Coulomb energy $\Gamma[u(\Gamma,N)-u_{0}]$ [54]. It is well known that, in contrast to the bulk system, phase transitions in finite-size systems are very diffuse, both in temperature and phase composition. As can be seen in Fig. 4, the BOCP is no exception. In this case, melting can be associated with the Coulomb coupling parameter $\Gamma_{\mathrm{m}}$ corresponding to the maximum of $\Gamma_{\mathrm{m}}[u(\Gamma_{\mathrm{m}},N)-u_{0}]$. This maximum is shifted toward lower values with the increase in NN. The inset shows a fit of $\Gamma_{\mathrm{m}}$ to the exponential decay function $\Gamma_{\mathrm{m}}(N)=174.0+647.995\exp(-N/1729.886)$, from which the estimate for the bulk melting parameter is $\Gamma_{\mathrm{m}}(\infty)=174\pm 2$. The error of 2 corresponds to the error of maximum resolution for the size-dependent function $\Gamma_{\mathrm{m}}[u(\Gamma_{\mathrm{m}},N)-u_{0}]$ due to the error in $u(\Gamma,N)$. Obtained value correlates with that obtained in [54] ($\Gamma_{\mathrm{m}}=178$) and coincides with the results by Dubin and O’Neil [17] and Khrapak and Khrapak [50]. The absence of an appreciable metastable region in the neighborhood of $\Gamma_{\mathrm{m}}$ is noteworthy. It means that the fluid–solid transition is very fast, at least, its characteristic time is less than $\tau N_{\tau}$. This can be connected with the solid-like shells near the BOCP boundary, which initiate crystallization propagating to the BOCP center.
@@
Figure 5: Thermal fraction of the electrostatic energy in the thermodynamic limit as a function of $\Gamma$. This work: MD results (dots) and their approximation by the function (38) (dashed line) and (39) with three fitting parameters (dashed-dotted line). Solid line indicates calculations based on the harmonic lattice model supplemented by anharmonicity correction [34] and dotted line presents MC results [54].
@@
Another way to estimate $\Gamma_{\mathrm{m}}$ is based on the parameter $\Gamma(u_{\infty}-u_{0})$ for the bulk fluid and solid (Fig. 5). Here, in contrast to Fig. 4, the dependence of this parameter on $\Gamma$ forms a cliff at $\Gamma_{\mathrm{m}}=174\pm 2$, where again, the estimation for accuracy follows from the error in $u_{\infty}(\Gamma)$. Note that at $190\leq\Gamma\leq 210$, the dots are situated well below the data for $\Gamma\geq 300$. This is a consequence of a poor applicability of the principal size dependence (19) for $N>10000$, owing to which the data in Table 2 were obtained only for the smaller sizes. Very large errors of $u_{\infty}$ calculations in this range of $\Gamma$ are indicative of enormous fluctuations of the BOCP structure at $\Gamma\approx\Gamma_{\mathrm{m}}$. Here, $N_{\tau}$ limited by the computational resources and listed in Table 1 seems to be insufficient to reach equilibration for such sizes. However, for $\Gamma\leq 178$ and $\Gamma\geq 300$, the systems are sufficiently close to equilibrium. One can testify that at high $\Gamma$, our data agree qualitatively with the analytical results by Chugunov and Baiko [34]. Note that, as for a finite size BOCP, we have found no metastable region. In addition, it is worth mentioning that the excess free energy at the melting point (28) amounts to $0.0092$, which proves to be an order of magnitude less than what follows from the estimate of free energy in [54].
@@
Figure 6: Size dependences of the excess interatomic electrostatic energy $u_{p{\mathrm{e}x}}$ (circles) and the excess ion–background electrostatic energy $u_{b{\mathrm{e}x}}$ (triangles) at $\Gamma=30$.
@@
Now we turn to the determination of the excess interatomic and ion–background electrostatic energy (Table 2 and Fig. 6). The size dependences of excess energies $u_{p\mathrm{e}x}(\Gamma,N)$ and $u_{b\mathrm{e}x}(\Gamma,N)$ (5) determined in our simulation reveal an overall trend to decrease with the increase in NN. However, the decrease is so small that we failed to assign any functional dependence to it in the investigated range of system parameters. Rigorously speaking, we are not able to extrapolate the results to the thermodynamic limit due to an extremely narrow abscissa range in Fig. 6 and hence, confine ourselves to its estimate by taking these quantities for $N=21200$. Here, the corresponding errors were calculated in the same way as for $u(\Gamma,N)$. The resulting dependences $u_{p{\mathrm{e}x}}(\Gamma)$ and $u_{b{\mathrm{e}x}}(\Gamma)$ are shown in Fig. 7 and Table 2. Since at $182\leq\Gamma\leq 210$, BOCP of the sizes $N>10000$ were not simulated, the corresponding $u_{p{\mathrm{e}x}}$ and $u_{b{\mathrm{e}x}}$ are not listed in the table. Behavior of these dependences is in agreement with the asymptotic analytical predictions (12) and (18), which is indicative of the fact that Fig. 7 represents a satisfactory estimate for the investigated quantities in the thermodynamic limit. One can see that as $\Gamma$ is decreased, $u_{b{\mathrm{e}x}}(\Gamma)$ vanishes more rapidly than $u_{p{\mathrm{e}x}}(\Gamma)$, which can be explained by the difference in $f_{p}(r)$ (7) and $f_{c}(r)$ (8). Indeed, at $\Gamma=0.1$, $f_{c}\simeq 1$ (Fig. 1a), which vanishes the integral (10), whereas $f_{p}$ is noticeably different from zero at short distances (Fig. 1b) resulting in nonzero $u_{p{\mathrm{e}x}}(\Gamma)$ (9). Since (9) and (10) were written in the approximation of pair correlations for a homogeneous medium, it is clear that such approximation is ineffective for $u_{b{\mathrm{e}x}}(\Gamma)$, whose analytical calculation would require the inclusion of higher-order correlations. In this connection, a fast decrease of $u_{b{\mathrm{e}x}}(\Gamma)$ at $\Gamma>0.3$ is a consequence of inadequacy of the Debye–Hückel approximation in this range of $\Gamma$. Notably, it follows from the recent analysis [66] of the reduced coefficients of self-diffusion, shear viscosity, and thermal conductivity of the OCP based on the Debye–Hückel plus hole approach [62] that the binary collision approaches break down at $\Gamma>0.3$, when OCP is no more a weakly coupled system.
@@
At high $\Gamma$, $u_{p{\mathrm{e}x}}(\Gamma)$ reaches a solid state value already at $\Gamma=30$ while $u_{b{\mathrm{e}x}}(\Gamma)$, only at $\Gamma=650$. Interestingly, $u_{p{\mathrm{e}x}}(\Gamma)$ has a shallow minimum at $\Gamma\approx 10$, which coincides with a gas-to-liquid dynamical crossover in the OCP discussed in works [29] and [65], so this minimum can serve as yet another definition of the Frenkel line in OCP.